OpenStudy (anonymous):

What is the common ratio of the following sequence? 2, -4, 8, -16 -1 2 1 -2

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

hey!!

OpenStudy (misty1212):

divide any term by the previous one that is all

OpenStudy (anonymous):

@misty1212 ^^

OpenStudy (anonymous):

ditto

OpenStudy (anonymous):

1?

OpenStudy (misty1212):

you can do \(\frac{-4}{2}\) or \(\frac{8}{-4}\) or \(\frac{-16}{8}\) makes no difference

OpenStudy (misty1212):

lol each of those numbers is the same, none of them are 1

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

verify -2

OpenStudy (anonymous):

thanks everyone